OpenStudy (here_to_help15):

Please help I will award medal and fan! Find the equation of the line that passes through the points (-1 , -1) and (-1 , 2).

OpenStudy (here_to_help15):

@camerondoherty

OpenStudy (here_to_help15):

@MLeighW

OpenStudy (anonymous):

find the slope which is given as y2-y1/x2-x1

OpenStudy (here_to_help15):

I get is this correct? To find the equation of the line through the points (-1 , -1) and (-1 , 2), we first use the slope m. m = (y2 - y1) / (x2 - x1) = (2 - -1) / (-1 - -1) = 3 / 0 The slope is indefined which means the line is perpendicular to the x axis and its equation has the form x = constant. Since both points have equal x coordinates -1, the equation is given by: x = -1

OpenStudy (here_to_help15):

Thanks anyways = )

OpenStudy (anonymous):

3/0 isn't possible so I have no idea :/